# A Unified Dynamical Field Theory of Learning, Inference, and Emergence

### v1

@@ Abstract
Learning, inference, and emergence in biological and artificial systems are often studied within disparate theoretical frameworks, ranging from energy-based models to recurrent and attention-based architectures. Here we develop a unified dynamical field theory in which learning and inference are governed by a minimal stochastic dynamical equation admitting a Martin–Siggia–Rose–Janssen–de Dominicis formulation. Within this framework, inference corresponds to saddle-point trajectories of the associated action, while fluctuation-induced loop corrections render collective modes dynamically emergent and generate nontrivial dynamical time scales. A central result of this work is that cognitive function is controlled not by microscopic units or precise activity patterns, but by the collective organization of dynamical time scales. We introduce the time-scale density of states (TDOS) as a compact diagnostic that characterizes the distribution of collective relaxation modes governing inference dynamics. Learning and homeostatic regulation are naturally interpreted as processes that reshape the TDOS, selectively generating slow collective modes that support stable inference, memory, and context-dependent computation despite stochasticity and structural irregularity. This framework unifies energy-based models, recurrent neural networks, transformer architectures, and biologically motivated homeostatic dynamics within a single physical description, and provides a principled route toward understanding cognition as an emergent dynamical phenomenon.
@@
Motivated by this analogy, we propose that cognition should be understood not as an explicit algorithmic computation, but as a collective dynamical phenomenon governed by an effective field theory. In this view, learning reshapes the dynamical landscape, inference corresponds to stable trajectories within that landscape, and emergent cognitive phenomena arise from the interplay between deterministic flow and stochastic fluctuations 27 ; 28 .
In this work, we introduce a unified dynamical field theory that describes learning, inference, and emergence within a single stochastic framework. Our starting point is a minimal dynamical equation for high-dimensional collective neural states, incorporating an effective potential shaped by learning, a state-dependent geometric structure, and non-conservative reentrant dynamics 29 . This formulation explicitly allows for recursive feedback and self-referential flows, which we argue are essential for stable inference and metacognitive behavior. We elevate this dynamical description to a Martin–Siggia–Rose–Janssen–de Dominicis (MSRJD) path-integral representation, enabling a systematic treatment of both average dynamics and fluctuations.
@@
Importantly, the framework advanced here entails a shift in what constitutes the fundamental object of inference. Rather than attributing cognitive function to individual neurons, synapses, or static activity patterns, we argue that inference is governed by collective dynamical modes distinguished by their characteristic time scales. In high-dimensional recurrent systems, spatial organization is largely fixed, while learning and adaptation act primarily by reshaping the temporal relaxation structure of the dynamics. We formalize this idea by introducing a time-scale density of states (TDOS), which quantifies the distribution of collective modes as a function of their relaxation rates. Learning reorganizes this distribution, selectively generating slow collective modes that support stable inference, memory, and context-dependent computation despite microscopic stochasticity, random connectivity, and structural irregularity.
The remainder of this paper is organized as follows. Section II introduces the unified dynamical equation governing collective neural states, and presents the corresponding MSRJD field-theoretic formulation. In Section III, we show that inference emerges as saddle-point dynamics, while Section IV demonstrates how loop corrections generate emergent collective modes and time scales. Section V describes learning as structural adaptation based on this concept, and Section VI derives several canonical learning architectures as special limits of the theory. Finally, Section VII discusses the broader implications of this framework for neuroscience and artificial intelligence and outlines directions for future research.
@@
representing either population neural activity in biological networks or latent representations in artificial systems. The collective state variable $x$ is a dimensionless coordinate in a high-dimensional state space, representing macroscopic neural activity. We propose that learning, inference, and emergent behavior are governed by the following minimal stochastic dynamical equation:
@@
Equation (1) constitutes a dynamical law for cognitive systems, in the sense that it provides a minimal and closed description of collective dynamics. The equation is not introduced heuristically; rather, it builds on the standard lineage from stochastic Langevin dynamics to variational path-probability formulations and geometric descriptions of diffusion on curved state spaces, and is here specialized to learning-shaped potentials, adaptive state-space geometry, and nonconservative reentrant flows. Within this framework, inference corresponds to relaxation along stable directions of the learned potential $\Phi$, while reentrant dynamics $R$ enable iterative reinterpretation of internal states beyond simple attractor dynamics. Figure 1(a) illustrates the conceptual structure of the unified dynamical field theory.
Equation (1) also admits a MSRJD path-integral representation. Introducing response fields $\tilde{x}(t)$, the generating functional reads
@@
This formulation provides a unified description of average inference trajectories (saddle points of $S$) and fluctuation-induced phenomena (loop corrections). In particular, critical slowing down and emergent collective modes arise naturally from infrared divergences of loop diagrams near soft directions of $\nabla^{2}\Phi$.
@@
Within the MSRJD formulation, inference corresponds to the most probable trajectory of the collective state conditioned on internal dynamics and stochastic fluctuations. This trajectory is obtained as the saddle point of the action $S[x,\tilde{x}$], and provides a dynamical realization of inference without invoking an explicit optimization algorithm.
@@
Equation (5) defines the mean inference trajectory of the system, representing the most probable evolution of the collective state in the weak-noise or large-$N$ limit. Importantly, inference is not formulated here as an explicit minimization or symbolic computation. Instead, it emerges as a relaxation flow on a learned dynamical landscape shaped jointly by the effective potential $\Phi(x)$, the state-space geometry encoded in $G(x)$, and reentrant dynamics $R(x)$, as illustrated in Fig. 1(b).
@@
Unlike static energy-based models, the inferred object here is not a single state but an entire temporal path. Inference is therefore inherently dynamical, encoding both the evolution of internal representations and their temporal consistency.
While geometric formulations of stochastic diffusion on curved state spaces have been extensively studied 41 ; 42 , the role of state-space curvature in the present framework is not to define manifold-valued noise, but to regulate the sensitivity and branching structure of deterministic inference trajectories. Beyond determining the mean inference trajectory, the state-space geometry induces an independent, purely geometric source of sensitivity. In particular, when the metric $G(x)$ is non-Euclidean, nearby saddle-point inference trajectories need not remain parallel. Instead, they experience relative acceleration due to the intrinsic Riemann curvature of the state space, even in the absence of noise or external perturbations. Depending on the local sectional curvature, this geometric effect can lead to either divergence (defocusing) or convergence (focusing) of neighboring trajectories. This phenomenon is directly analogous to tidal forces in general relativity, where spacetime curvature causes freely falling geodesics to separate or focus without the action of external forces.
Formally, the separation vector $\zeta^{i}(t)$ between two nearby inference trajectories obeys a geodesic deviation equation of the form
@@
where $R^{i}{}_{jkl}[G$] is the Riemann curvature tensor associated with the state-space metric. We interpret this curvature-induced relative acceleration as a cognitive tidal force: a purely geometric mechanism that regulates inference sensitivity by inducing either branching or focusing of trajectories, independently of stochastic fluctuations (See the details in Appendix C).
@@
The reentrant term $R(x)$ plays a crucial role in this process. When $R=0$, the dynamics reduces to pure gradient descent, leading to rapid convergence to fixed points determined solely by $\Phi$. While sufficient for static memory retrieval, such dynamics lacks the capacity for iterative reinterpretation and contextual modulation. Non-conservative reentrant flows introduce rotational components in state space, enabling trajectories to explore learned manifolds before settling and thereby supporting recursive computation and context-dependent inference.
@@
The saddle-point description captures only the leading-order behavior. As shown in the next section, fluctuations around these trajectories generate additional collective modes and emergent time scales through loop corrections, providing a field-theoretic mechanism for emergence and criticality.
@@
IV.2 B. Loop Corrections, Renormalization, and Time-Scale Density of States
@@
IV.3 C. Emergent Collective Modes and Physical Interpretation
@@
defines a dressed collective relaxation mode with lifetime $\tau\sim m^{-1}$. While the corresponding slow mode exists already at the bare (mean-field) level with rate $m_{0}$, fluctuation-induced self-energy effects renormalize and regulate it, producing a self-generated infrared scale even when $m_{0}\to 0$. We interpret such regulated quasi-soft excitations as carriers of emergent computation and memory, analogous to quasiparticles in condensed-matter systems. We define a neuroton as a self-generated collective relaxation mode arising from fluctuation-induced dressing around a reentrant saddle-point dynamics.
More generally, the loop expansion reveals a universal mechanism underlying emergence: whenever nonlinear interactions and stochastic fluctuations coexist, the system necessarily generates new collective time scales beyond mean-field dynamics. Within the present framework, saddle-point dynamics implements inference, loop corrections generate emergent collective modes, and critical behavior arises from the competition between deterministic drift and fluctuation-induced self-energy.
@@
In the present theory, learning is not an independent algorithmic procedure, nor an external optimization rule imposed on the dynamics. Instead, learning is identified with a slow structural adaptation of the unified dynamical system, driven by the statistics of inference trajectories and their associated collective modes. Crucially, learning acts by reorganizing the distribution of dynamical time scales uncovered by the loop analysis, thereby shaping the temporal architecture of cognition.
@@ A. Learning as Slow Structural Adaptation
Equation (22) expresses the principle that learning is driven by the accumulated statistics of inference rather than by instantaneous fluctuations. Repeated inference episodes under similar inputs bias the structural parameters so as to stabilize frequently visited trajectories and destabilize inconsistent ones. Learning therefore sculpts the geometry of state space, making correct inference dynamically natural rather than algorithmically enforced.
@@ B. Learning as Reorganization of the Time-Scale Density of States
From this perspective, learning operates by reorganizing the time-scale density of states associated with collective modes. Structural adaptation selectively enhances slow collective modes associated with behaviorally relevant features of the input statistics, while suppressing fast, unstable fluctuations. Importantly, a single collective mode does not encode information; rather, cognition emerges from the structured distribution of modes across time scales.
@@
From the present perspective, RNNs correspond to discrete-time approximations of the unified continuous dynamics, in which inference is implemented as iterative relaxation under a locally linearized effective potential. The recurrence in RNNs thus provides a rudimentary form of feedback, while the discrete time step $\Delta t$ imposes an externally fixed temporal structure.
Element-wise nonlinearities such as sigmoidal activation functions introduce local, state-dependent gain modulation and play an important role in stabilizing the dynamics. However, these nonlinearities act at the level of individual units and do not modify the geometry of the collective state space itself. As a result, while RNNs can represent temporal dependencies and sequential structure, they lack mechanisms for learning state-space geometry or dynamically reshaping the distribution of collective time scales.
@@
Within the unified dynamical field theory, this recurrence corresponds to a fixed reentrant-like flow acting on an augmented state space. The normalization by $Z_{t}$ induces an implicit, architecture-prescribed scaling that may be interpreted as a constrained or fixed metric structure. Crucially, however, both the recurrence rule and the effective geometry are externally specified by the model design and do not evolve dynamically with the system state.
@@
Consequently, transformers realize a discrete-time recurrent approximation of the unified dynamical equation, enabling efficient integration of long-range dependencies through structured memory accumulation. At the same time, they lack mechanisms for state-dependent reorganization of collective mode structure or for dynamically reshaping the distribution of intrinsic time scales. In this sense, transformer architectures occupy an intermediate regime: they go beyond purely feedforward models by implementing recurrence at the level of collective memory, yet remain distinct from genuinely reentrant systems in which the effective geometry and stability structure are dynamically modulated by the system’s own activity.
@@
Equation (37) stabilizes activity on a finite-radius shell for $\gamma\rho>1$, strongly suppressing radial fluctuations. This stabilization can be interpreted as the emergence of an anisotropic metric on state space, in which radial directions are stiff while tangential directions remain soft. Figure 3 provides explicit visualizations of the effective homeostatic potential emerging from the FHRN dynamics. These figures complement the analytical derivation by illustrating how radial stabilization and angular degeneracy jointly produce a shell-shaped attractor manifold.
@@
Figure 4 displays the TDOS of the FHRN, extracted from the spectrum of the linearized dynamics around a stable homeostatic fixed point, revealing a dense accumulation of slow collective modes. The pronounced accumulation of modes at small relaxation rates $\lambda$ reveals a dense spectrum of slow collective modes generated by the interplay of homeostatic stabilization and antisymmetric reentrant mixing. These long-lived modes dominate inference dynamics and provide a dynamical substrate for memory and temporal integration. The continuous structure of the TDOS highlights that cognition in the FHRN is governed by the collective organization of time scales rather than by isolated modes or fixed points.
@@
A central conceptual implication of this theory is a reinterpretation of inference as a dynamical process rather than an algorithmic operation. In the present framework, inference is realized as relaxation along stable directions of a learned potential landscape embedded in a nontrivial state-space geometry. Learning does not compute solutions explicitly; instead, it sculpts the effective potential $\Phi$ and the metric $G$, thereby shaping the geometry on which inference unfolds. This geometric perspective naturally explains the robustness of inference across diverse implementations, from biological neural circuits to artificial architectures, despite substantial microscopic variability and noise.

### v2

@@ Abstract
Learning, inference, and emergence in biological and artificial systems are often studied within disparate theoretical frameworks, ranging from energy-based models to recurrent and attention-based architectures. Here we develop a unified dynamical field theory in which learning and inference are governed by a minimal stochastic dynamical equation admitting a Martin–Siggia–Rose–Janssen–de Dominicis formulation. Within this framework, inference corresponds to saddle-point trajectories of the associated action, while fluctuation-induced loop corrections render collective modes dynamically emergent and generate nontrivial dynamical time scales. A central result of this work is that cognitive function is controlled not by microscopic units or precise activity patterns, but by the collective organization of dynamical time scales. We introduce the time-scale density of states (TDOS) as a compact diagnostic of the distribution of collective relaxation modes governing inference dynamics. Learning and homeostatic regulation are naturally interpreted as processes that reshape both the effective potential and the underlying state-space geometry, thereby reorganizing the TDOS and selectively stabilizing slow collective modes that support stable inference, memory, and context-dependent computation despite stochasticity and structural irregularity. This framework unifies energy-based models, recurrent neural networks, transformer architectures, and biologically motivated homeostatic dynamics within a single physical description, and provides a principled route toward understanding cognition as an emergent dynamical phenomenon.
@@
Motivated by this analogy, we propose that cognition should be understood not as an explicit algorithmic computation, but as a collective dynamical phenomenon governed by an effective field-theoretic description of state-space dynamics. In this view, learning reshapes the dynamical landscape, inference corresponds to stable trajectories within that landscape, and emergent cognitive phenomena arise from the interplay between deterministic flow and stochastic fluctuations 27 ; 28 .
In this work, we introduce a unified dynamical field theory that describes learning, inference, and emergence within a single stochastic framework. Our starting point is a minimal dynamical equation for high-dimensional collective neural states, incorporating an effective potential shaped by learning, a state-dependent geometric structure that governs sensitivity and stability, and non-conservative reentrant dynamics 29 . This formulation explicitly allows for recursive feedback and self-referential flows, which we argue are essential for stable inference and metacognitive behavior. We elevate this dynamical description to a Martin–Siggia–Rose–Janssen–de Dominicis (MSRJD) path-integral representation, enabling a systematic treatment of both average dynamics and fluctuations.
@@
Importantly, the framework advanced here entails a shift in what constitutes the fundamental object of inference. Rather than attributing cognitive function to individual neurons, synapses, or static activity patterns, we argue that inference is governed by collective dynamical modes distinguished by their characteristic time scales. In high-dimensional recurrent systems, spatial organization is largely fixed, while learning and adaptation act primarily by reshaping the temporal relaxation structure of the dynamics. We formalize this idea by introducing a time-scale density of states (TDOS), which quantifies the distribution of collective modes as a function of their relaxation rates. Learning reorganizes this distribution through the joint modulation of dynamical stability and state-space geometry, selectively generating slow collective modes that support stable inference, memory, and context-dependent computation despite microscopic stochasticity, random connectivity, and structural irregularity.
The remainder of this paper is organized as follows. Section II introduces the unified dynamical equation governing collective neural states, and presents the corresponding MSRJD field-theoretic formulation. In Section III, we show that inference emerges as saddle-point dynamics, while Section IV demonstrates how loop corrections generate emergent collective modes and time scales. Section V describes learning as structural adaptation, and Section VI derives several canonical learning architectures as special limits of the theory. Finally, Section VII discusses the broader implications of this framework for neuroscience and artificial intelligence and outlines directions for future research.
@@
representing macroscopic neural activity in biological networks or latent representations in artificial systems. The state variable $x$ is treated as a dimensionless coordinate in a high-dimensional state space, capturing collective degrees of freedom rather than individual microscopic units.
We propose that learning, inference, and emergent behavior are governed by the following minimal stochastic dynamical equation:
@@
Equation (1) constitutes a dynamical law for cognitive systems in the sense that it provides a minimal and closed description of collective dynamics. Rather than being introduced heuristically, this equation follows the standard lineage from stochastic Langevin dynamics to variational path-probability formulations. Within this framework, inference corresponds to relaxation along stable directions of the learned potential $\Phi(x)$, while the reentrant flow $R(x)$ enables iterative reinterpretation of internal states beyond simple attractor dynamics. Figure 1 schematically illustrates the conceptual structure of the unified dynamical field theory.
II.2 B. Path-Integral Formulation and Emergent Collective Modes
Equation (1) admits a MSRJD path-integral representation. Introducing response fields $\tilde{x}(t)$, the generating functional is given by
@@
This formulation provides a unified description of both average inference trajectories, corresponding to saddle points of the action, and fluctuation-induced phenomena captured by loop corrections. In particular, critical slowing down and emergent collective modes arise naturally from infrared divergences of loop diagrams near soft directions of the Hessian $\nabla^{2}\Phi$. The MSRJD formalism thus establishes a direct link between stochastic dynamics, inference, and emergence within a single theoretical framework.
II.3 C. Geometric Origin of the Unified Dynamical Equation
Beyond its stochastic and variational motivation, the unified dynamical equation admits a direct geometric interpretation when formulated on a Riemannian state space. We begin by viewing the collective state $x\in\mathbb{R}^{N}$ as a coordinate chart on a differentiable manifold endowed with a generally non-Euclidean metric $G(x)$. When the metric is state dependent, the induced geometry is intrinsically Riemannian, even though the coordinates themselves take values in $\mathbb{R}^{N}$.
On a Riemannian manifold, the most general form of a gradient flow generated by a scalar potential $\Phi(x)$ is given by
$\dot{x}^{i}=-G^{ij}(x)\,\partial_{j}\Phi(x),$ (5)
where $G^{ij}(x)$ denotes the inverse metric. Equation (5) represents the canonical notion of steepest descent with respect to the geometry defined by $G(x)$, and reduces to ordinary Euclidean gradient descent only when the metric is flat and constant.
However, a general vector field on a Riemannian manifold is not exhausted by gradient flows alone. Given a specified metric, any sufficiently regular drift field admits a natural decomposition into a metric-dependent gradient component and an independent, non-conservative component that is orthogonal, with respect to the inner product induced by $G(x)$, to all metric gradient vector fields. Allowing for this geometrically unavoidable remainder leads to the deterministic dynamics
$\dot{x}^{i}=-G^{ij}(x)\,\partial_{j}\Phi(x)+R^{i}(x),$ (6)
where $R(x)$ denotes a solenoidal flow that cannot be expressed as the gradient of any scalar functional.
This decomposition is closely related to Helmholtz–Hodge–type decompositions of vector fields, suitably generalized to curved (Riemannian) state spaces under standard regularity conditions 41 . Within this interpretation, the potential $\Phi(x)$ selects preferred directions of relaxation, the metric $G(x)$ determines how sensitivity is distributed across the state space, and the solenoidal term $R(x)$ captures autonomous circulation along dynamically stabilized manifolds.
This decomposition highlights a qualitative distinction between gradient-driven relaxation and reentrant dynamics. The gradient term implements inference as relaxation toward stable manifolds shaped by learning, whereas the non-conservative flow $R(x)$ redistributes activity along these manifolds without altering the value of the effective potential. Such circulation enables iterative reinterpretation, contextual modulation, and temporal looping of internal representations, extending beyond simple attractor dynamics.
Incorporating a geometrically consistent reentrant flow $R(x)$ constitutes a principled completion of the dynamical structure suggested by the underlying geometry. The unified dynamical equation thus represents the most general geometrically consistent form of a cognitive field on a Riemannian state space, providing a closed description that simultaneously accommodates learning-shaped potentials, adaptive state-space geometry, and non-conservative reentrant dynamics.
@@
Within the MSRJD formulation, inference corresponds to the most probable trajectory of the collective state conditioned on internal dynamics and stochastic fluctuations. This trajectory is obtained as the saddle point of the action $S[x,\tilde{x}$], providing a dynamical realization of inference without invoking an explicit optimization algorithm.
@@
Equation (7) defines the mean inference trajectory of the system, representing the most probable evolution of the collective state in the weak-noise or large-$N$ limit. Inference is not formulated here as an explicit minimization or symbolic computation. Instead, it emerges as a relaxation flow on a learned dynamical landscape shaped jointly by the effective potential $\Phi(x)$, the state-space geometry encoded in $G(x)$, and reentrant dynamics $R(x)$, as illustrated in Fig. 1(b).
@@
Unlike static energy-based models, the inferred object here is not a single state but an entire temporal path. Inference is therefore inherently dynamical, encoding both the evolution of internal representations and their temporal consistency. In the present framework, the saddle-point trajectory provides the deterministic backbone along which slower collective modes, generated by fluctuations, may subsequently emerge and persist.
While geometric formulations on curved state spaces have been explained in Section II, the role of state-space curvature in the present context is to regulate the sensitivity and branching of deterministic inference trajectories. Here, the metric $G(x)$ enters not as an abstract geometric object, but as an effective Fisher information metric induced by the local probabilistic sensitivity of inference dynamics 42 .
When the metric $G(x)$ is non-Euclidean, nearby saddle-point inference trajectories need not remain parallel, even in the absence of noise or external perturbations. Instead, they experience relative acceleration due to the intrinsic Riemann curvature of the state space. Depending on the local sectional curvature, this geometric effect can lead to either divergence (defocusing) or convergence (focusing) of neighboring trajectories, in direct analogy with tidal effects in general relativity.
Formally, the separation vector $\zeta^{i}(t)$ between two nearby inference trajectories obeys the geodesic deviation equation
@@
where $R^{i}{}_{jkl}[G$] is the Riemann curvature tensor associated with the state-space metric. We interpret this curvature-induced relative acceleration as a cognitive tidal force: a geometric manifestation of the information-theoretic sensitivity encoded in the Fisher metric, regulating the branching or focusing of inference trajectories independently of stochastic fluctuations (see Appendix C).
@@
The reentrant term $R(x)$ plays a crucial role in this process. When $R=0$, the dynamics reduces to pure gradient descent, leading to rapid convergence to fixed points determined solely by $\Phi$. While sufficient for static memory retrieval, such dynamics lacks the capacity for iterative reinterpretation and contextual modulation. Beyond exploration, such reentrant flows play a crucial role in preserving tangential degrees of freedom along stabilized manifolds, thereby enabling the emergence of long-lived collective modes discussed in Sec. IV.
@@
The saddle-point description captures only the leading-order, deterministic backbone of inference. As shown in the next section, fluctuations around these trajectories generate additional collective modes and emergent time scales through loop corrections, providing a field-theoretic mechanism for emergence and criticality.
@@
Although the reentrant flow $R(x)$ does not primarily set the dissipative relaxation rates at the Gaussian level, its gradient generically induces antisymmetric and non-normal components in the linearized stability matrix. These components generate circulation and transient amplification in state space, modifying phase structure, trajectory geometry, and sensitivity, while the dominant asymptotic relaxation rates remain controlled by the symmetric, potential-derived contribution $G^{-1}\nabla^{2}\Phi$.
IV.2 B. Propagators and Collective Modes in Joint State Space
Before interpreting the loop-corrected response functions, it is important to clarify the nature of the degrees of freedom being propagated. Although the unified dynamical equation is written in terms of a state vector $x(t)$, the retarded propagator $G_{R}$ does not transmit states themselves. Rather, it governs the evolution of infinitesimal perturbations $\delta x(t)$ in the tangent space around an inference trajectory $x^{*}(t)$.
In sequence-processing systems such as recurrent networks or transformers, the state vector $x$ naturally represents a joint configuration of multiple tokens, $x=(x_{1},\dots,x_{T})\in\mathbb{R}^{T\times d}$. Linearization around a saddle-point trajectory therefore defines a stability matrix $M$ acting on the joint token state space. Its eigenvectors correspond to collective directions spanning multiple tokens and representation dimensions, while its eigenvalues set the characteristic relaxation rates of these collective modes.
The propagator of Eq. (15) thus decomposes naturally into a sum over collective eigenmodes,
$G_{R}(\omega)=\sum_{\alpha}\frac{|\psi_{\alpha}\rangle\langle\tilde{\psi}_{\alpha}|}{-i\omega+\lambda_{\alpha}},$ (17)
where $\psi_{\alpha}$ denote collective modes in the joint state space. Apparent token-to-token influence arises from the projection of these collective modes back onto individual token coordinates. In this sense, propagation in the unified dynamical theory is mediated not by individual states, but by collective modes spanning the full state-space configuration.
More precisely, the collective modes governing fluctuation propagation are defined as eigenmodes of the linearized stability matrix $M$ acting on the joint state space,
$M\,\psi_{\alpha}=\lambda_{\alpha}\,\psi_{\alpha},$ (18)
where $\psi_{\alpha}$ are eigenvectors spanning coordinated directions across multiple tokens and representation dimensions, and $\lambda_{\alpha}$ denote the associated relaxation rates.
In this formulation, a single eigenmode $\psi_{\alpha}$ generally does not correspond to a localized perturbation of an individual token, but to a distributed collective pattern in the full token configuration space. Slow collective modes arise when one or more eigenvalues $\lambda_{\alpha}$ approach zero, signaling the emergence of long-lived dynamical structures.
IV.3 C. Loop Corrections, Renormalization, and Time-Scale Density of States
@@
IV.4 D. Emergent Collective Modes and Physical Interpretation
@@
defines a dressed collective relaxation mode with lifetime $\tau\sim m^{-1}$. While such slow modes exist already at the bare (mean-field) level with rate $m_{0}$, fluctuation-induced self-energy effects renormalize and regulate it, generating a self-organized infrared scale even in the limit $m_{0}\to 0$.
We interpret such regulated quasi-soft excitations as carriers of emergent computation and memory, analogous to quasiparticles in condensed-matter systems. Specifically, we define a neuroton as a self-generated collective relaxation mode arising from fluctuation-induced dressing around a reentrant saddle-point dynamics. Neurotons correspond to long-lived collective excitations associated with low-frequency poles of the retarded Green’s function in the joint state space, representing coordinated dynamical patterns spanning multiple tokens and representation dimensions near the homeostatically stabilized shell. In this sense, they are analogous to Goldstone-like modes, representing protected angular fluctuations whose effective energy scale is set by inverse relaxation times rather than microscopic Hamiltonians.
At a more general level, the loop expansion reveals a universal mechanism for emergence: whenever nonlinear interactions and stochastic fluctuations coexist, new collective time scales are dynamically generated beyond mean-field dynamics. Within the present framework, saddle-point dynamics implements inference, loop corrections give rise to emergent collective modes such as neurotons, and critical behavior arises from the competition between deterministic drift and fluctuation-induced self-energy.
@@
In the present theory, learning is not an independent algorithmic procedure, nor an external optimization rule imposed on the dynamics. Instead, learning is identified with a slow structural adaptation of the unified dynamical system, driven by the statistics of inference trajectories and their associated collective modes. Crucially, learning acts by reorganizing the distribution of dynamical time scales uncovered by the loop analysis, thereby shaping the temporal architecture of cognition. Equivalently, learning reshapes the spectrum of collective modes—neurotons—that can be sustained by the dynamics, thereby controlling which long-lived trajectories remain available for inference.
@@ A. Learning as Slow Structural Adaptation
Equation (26) expresses the principle that learning is driven by the accumulated statistics of inference rather than by instantaneous fluctuations. Repeated inference episodes under similar inputs bias the structural parameters so as to stabilize frequently visited trajectories and destabilize inconsistent ones. From the perspective of Sec. IV, this stabilization corresponds to the selective reinforcement of slow collective modes in the joint state space, while fast-decaying modes are progressively suppressed. Learning therefore sculpts the geometry of state space, making correct inference dynamically natural rather than algorithmically enforced.
@@ B. Learning as Reorganization of the Time-Scale Density of States
From this perspective, learning operates by reorganizing the time-scale density of states associated with collective modes. Structural adaptation selectively enhances slow collective modes associated with behaviorally relevant features of the input statistics, while suppressing fast, unstable fluctuations. Individual collective modes (neurotons) provide the dynamical carriers of computation, but meaningful information is encoded only at the level of their organized temporal spectrum, as captured by the TDOS.
@@
From the present perspective, RNNs can be viewed as discrete-time approximations of the unified continuous dynamics, in which inference is implemented as iterative relaxation under a locally linearized effective potential. The recurrence in RNNs thus provides a basic form of temporal feedback, while the discrete time step $\Delta t$ imposes an externally fixed time scale.
More generally, if the effective recurrent interaction contains an antisymmetric component, the resulting discrete-time dynamics may also exhibit nonconservative circulation, corresponding to the reentrant term $R(x)$ in the continuous formulation. Such effects can give rise to non-normal dynamics and transient amplification beyond pure gradient relaxation, but they are not structurally constrained and need not be confined to stable manifolds. As a result, such circulation does not generically give rise to protected long-lived collective modes or a reorganized time-scale density of states.
Element-wise nonlinearities such as sigmoidal activation functions introduce local, state-dependent gain modulation and play an important role in stabilizing individual units. However, these nonlinearities do not induce a learned geometry of the collective state space. As a result, while RNNs can represent temporal dependencies and sequential structure, they lack intrinsic mechanisms for dynamically shaping state-space geometry or redistributing collective time scales through self-generated collective modes.
@@
Within the unified dynamical field theory, this recurrence corresponds to a fixed reentrant-like flow acting on an augmented state space, with the qualifier that its structure is externally specified rather than dynamically learned. The normalization by $Z_{t}$ induces an implicit, architecture-prescribed scaling that may be interpreted as a constrained or fixed metric structure. Crucially, however, both the recurrence rule and the effective geometry are externally specified by the model design and do not evolve dynamically with the system state.
@@
Consequently, transformers realize a discrete-time recurrent approximation of the unified dynamical equation, enabling efficient integration of long-range dependencies through structured memory accumulation. At the same time, they lack mechanisms for state-dependent reorganization of collective mode structure or for dynamically reshaping the distribution of intrinsic time scales. As a result, long-lived collective excitations in transformers remain architecture-imposed rather than dynamically generated. In this sense, transformer architectures occupy an intermediate regime: they go beyond purely feedforward models by implementing recurrence at the level of collective memory, yet remain distinct from genuinely reentrant systems in which the effective geometry and stability structure are dynamically modulated by the system’s own activity.
@@
Equation (41) stabilizes activity on a finite-radius shell for $\gamma\rho>1$, strongly suppressing radial fluctuations. This stabilization can be interpreted as the emergence of an anisotropic metric on state space, in which radial directions are stiff while tangential directions remain soft. From an information-geometric perspective, this dynamically generated anisotropic metric admits a natural interpretation as an effective Fisher information metric induced by homeostatic suppression of radial uncertainty, as detailed in Appendix D. Figure 3 provides explicit visualizations of the effective homeostatic potential emerging from the FHRN dynamics. These figures complement the analytical derivation by illustrating how radial stabilization and angular degeneracy jointly produce a shell-shaped attractor manifold.
@@
Figure 4 displays the TDOS of the FHRN, extracted from the spectrum of the linearized dynamics around a stable homeostatic fixed point, revealing a dense accumulation of slow collective modes. The pronounced accumulation of modes at small relaxation rates $\lambda$ reveals a dense spectrum of slow collective modes generated by the interplay of homeostatic stabilization and antisymmetric reentrant mixing. These long-lived modes dominate inference dynamics and provide a dynamical substrate for memory and temporal integration. These modes correspond precisely to the neurotons introduced in Sec. IV. The continuous structure of the TDOS highlights that cognition in the FHRN is governed by the collective organization of time scales rather than by isolated modes or fixed points.
@@
VII.1 A. Inference, Geometry, and Emergence as a Unified Dynamical Principle
@@
A central conceptual implication of this theory is a reinterpretation of inference as a dynamical process rather than an algorithmic operation. In the present framework, inference is realized as relaxation along stable directions of a learned potential landscape embedded in a nontrivial state-space geometry. Learning does not compute solutions explicitly; instead, it sculpts the effective potential $\Phi$ and the metric $G$, thereby jointly shaping both the landscape and the geometry on which inference unfolds. This geometric perspective naturally explains the robustness of inference across diverse implementations, from biological neural circuits to artificial architectures, despite substantial microscopic variability and noise.